There are three major ferry lines operating on the East River, each generating 200 units of pollution per year. The ferry lines face the costs of reducing pollution that are shown in the table below.

Ferry line Cost of reducing pollution by 1 unit
A $1,500
B $2,400
C $1,800





a. The government has decided it wants to reduce pollution levels by 50 percent and requires each ferry line to cut its pollution in half. Determine the cost of pollution reduction for each ferry line.

Instructions: Round your answers to the nearest dollar.

Ferry line Cost of pollution reduction
A $
B $
C $


The total cost of pollution reduction is

    

.

b. Suppose the government decides to give each ferry line 100 tradable permits. Each permit allows a ferry line to produce one unit of pollution.

    Ferry line B  Correctwill buy  Not attemptedpermits from ferry line A  Correct.

c. Under this system of tradable permits, the total cost of pollution reduction is

please try to explain how to do it. I think I got it but if my answer is wrong I need to know how to do it! thank you!
Answer #1

Ans .

Given to us-

Generates 200 units of Pollution

Ferry Line Cost of reducing Pollution ( 1 Unit)
A $1500
B $2400
C $1800

a)

Reduce Pollution by 50% means 200 x 50% = 100 units

Ferry Line Cost of reducing Pollution ( 100 Unit)
A $1500x100 = $150,000
B $2400x100 = $240,000
C $1800x100 = $180,000

Total Cost of reduction = 150,000 + 240,000 + 180,000 = $570,000

b)

Because of ferry line B having the most cost.So Ferry line B will buy the tradable permit from ferry line A.

Ferry line B will buy permits 100 permits from ferry line A .

c)

Under the tradeable permits, the Cost of Reduction will be as follows.

Ferry Line Cost of reducing Pollution ( 100 Unit)
A $1500x200 = $300,000
B $2400x0= 0
C $1800x100 = $180,000

Under tradeable permits' cost of pollution reduction will be = 300,000 + 0 + 180,000 = $480,000

----------------------------------------------